Fermi National Laboratory
 

Accelerator Update for September 14, 2007 - September 21, 2007

 

Fermilab Shutdown Report
9/14/07 – 9/21/07

LINAC

Things Done This Week

  • Completed This Week
    • Adjusted Drive Loop on LRF5
    • Installed VXI Crates in LRF2, 3 & 1
    • Water System Maintenance
    • Shutoff All Systems in Preparation for Water Maintenance and Power Outage
    • LRF5 Cleaning
  • To do next week
    • Test Klystron Transformer Oil
    • Establish Baseline for HERF LLRF Calibration
    • Install VXI Crate in LRF4
    • Find Oscillation Problem in LRF4 Modulator
    • Repair LRF5 Modulator Weakness
    • Continue LLRF Studies in LRF5
  • The Neutron Therapy Facility is treating patients during the evening
    • After success of 9" line cavity coupler loop tuning on LRF3, we will repeat the procedure on LRF5 to reduce reverse during beam time. This will start on Tuesday, require beam pulses to verify beam loading / reverse power that night if all goes well, then return shielding blocks on Wednesday.
    • Start and Finish LLRF VXI Installation on LRF1 & LRF2
    • Install Co-axial cables needed for LLRF Controls
    • Install controls electronics in L.E. Linac


Booster

Water work:

  • LCW water work in the Booster West Gallery
    • Booster 95LCW was filled back up and hydrostatic tested
      • Few leaks were found and repaired
      • Tested BEG & BWG LCW pressure regulating valves
      • Increase pressure to nominal (110 psi) today
    • Most of the BRF stations at BWG are valved in
  • CUB
    • Power outage for CUB this weekend will prevent water to the system
      • Efforts continue Sunday after CUB turn pumps back on
  • Corrector installation
    • Number of new correctors in the tunnel: 12
      • L19, L17, L21, L23, L15, L9, L11, L13, L5, L7, L1, & L3
    • L3 corrector is fully installed
  • BPM field compensation box has been installed at 10 locations
  • Next week: Continue monitoring vacuum Complete the electronics installation
  • Additional Activities
    • Q800 replacement is complete
      • Currently recovering vacuum
    • Reinstallation of the 400MeV elements is complete
      • Currently recovering vacuum
  • Looking ahead
    • Coordinating final checks, walkthroughs and clean up in preparation for Oct. 3rd startup
      • Commissioning plans are under development

Antiproton Source

Mechanical

  • Leak Check (4)
    • Finished D/A, A40, A30, AP3
      • D:ESEP, A30 IP fixed; AP3 Pump
  • D:ISEP Shielding
  • Ion Chamber removal
  • Deb DCCT motorized stand
  • D:EKIK ceramic tube
    • Aligned; final connections to do
  • AP2 Water Hose Replacement
    • Progressing
  • Leaking LQA
    • Replaced; await water
  • Replaced leaking Debuncher Sextuple
  • SMA water leak
    • Surveyed into place
  • 4 water leaks to fix
    • IQ8 fixed twice
  • A10 Obstruction Search
    • Opened; components to shuffle

Target Station

  • Water Systems (2)
  • Temperature Monitoring (2)
    • 1 wiring done
  • Target Rotation
  • Storage
  • SEM Module
  • LVDTs
  • Sample Retrieval
  • Lens motion
  • Air Exhaust Stack
  • 10mm-2 into transformer
    • Prepping; await tools
  • 10mm-3 into transformer
    • Prepping, installing
  • Lens replacement
    • Started
  • Hipot of PMAG

Cryo

  • Valve Stem replace
  • Leak Check (1 of 2)
  • LN2 transfer line maintenance

Cooling Systems

  • Deb 3&4 Trans Notch Filters
  • Deb Double Notch Filters
  • Core 4-8GHz Equalizer
  • TWT balance & trip points
  • Stacktail Filter 3 cable change

Electrical (Pbar & EE Support)

  • Shunt Move
    • Progress water done
  • Shunt Repairs (9)
    • Nearly all rebuilt
  • PS Maintenance
    • Started
  • Lens SCRs
  • Regulation Electronics
    • Started by EE
  • Ion Pump maintenance
    • Started; await fans
  • Deb Shunt re-arrangement
  • Deb Extraction kicker load adjustment

Other

  • Interlocks/Safety System (3)
  • Repair cathodic protection
  • Holes under buildings
    • Water dripping from tunnel ceiling
      • Contractors sealed ceiling
  • DCCT calibration
    • Started
  • Baking System checkout
  • RF & TWT filters maintenance
  • Lights Checkout
  • Acc BPM repairs
    • Started
  • Activation Survey
  • Inductance checkout
  • PM ODH Fans & Sumps
  • Tunnel Wireless
  • Install AP2 Leak Detector
  • SEM maintenance
  • Clogged sump line – Transport
    • Berm dug; Replaced entire line
    • Needs Survey

Main Injector

  • MI40 beam tube replacement
    • Replacing beam pipes on the upstream end with larger aperture pipe.
    • Replacing downstream pipe due to alignment problems
  • MI30 collimators
    • Making vacuum connections
    • Buss and LCW moves mostly complete
    • Installing magnetic shielding
    • Installing marble and concrete masks
    • Installing and checking controls
  • MI Status

Recycler

  • Finished or underway
    • New DCCT section installed, leak checked and under vacuum since Wednesday last week; pumping down
      • DCCT section has been aligned
      • Beam pipes in this section (between valves) will be baked to 150 degrees C, the DCCT unit will be baked to 100 degrees C, starting today (Sept. 21st)
    • V303 moved, connected
    • H308 moved, connected
    • STC tank 7 leak check finished
      • Second attempt has been made; detected air but found no external leak
        • Sublimated TSPs at STCs 7,8,9,10.11,12
    • Sublimated TSPs from 204 to 210
    • Trombone additional quadrupoles installation issues
      • Installation has been completed
      • Measured rough positions of trombone elements with respect to surveyed elements
      • Connections started
    • Added support to old DCCT
    • Bulk supply upgrade (on-going)
    • DCCT cable connected
    • Stochastic cooling: Photodiodes aligned
    • The 800 transfer line is now leak free and being baked until this Friday (9/21).
      • MW803 seal has been replaced; leak fixed!
        • MW803 survey requested and has been done…(no sticker)
      • HP803 (large elliptical BPM) has been replaced with a standard elliptical unit
        • Leak free!
        • Survey of as-found has been done between gradient magnets
      • For the Recycler portion of the transfer line, there is only one large elliptical BPM is left (HP804).
      • Jim Fitzgerald is measuring the responses of spare Large elliptical BPMs
    • TOR330 repair under way (tunnel part is done)
    • BPM calibration has started (Peter Prieto)
    • QCL cable from 301 to 309 has been re-positioned; will be connected today
  • Recycler to do list
  • Magnet moves and survey /alignment pending calculations (we will have numbers next week)
  • Elevation survey after collimator installation
  • Field check for new quadrupoles
  • LEP dipole field centering on beampipes
  • General vacuum maintenance
  • Kicker repair; timing verification
  • Diagnostics maintenance
    • FW paddles
    • BLM for scrapers
    • Schottky (1.75 GHz) system gain calibration
    • Schottky (21 MHz) maintenance
  • Laser link vacuum maintenance
  • Damper status (in two weeks)
  • End-of-shutdown checklist ready (Cons)
  • Start-up plan in progress (Recycler)
  • Recycler to do list
    • Magnet moves and survey /alignment pending calculations
    • Elevation survey after collimator installation
    • Electrical connection for new quadrupoles
    • Field check for new quadrupoles
    • LEP dipole field centering on beampipes
    • Vacuum maintenance
    • Kicker repair; timing verification
    • TOR330 repair
    • BPM calibration
    • Diagnostics maintenance
    • Laser link vacuum maintenance

Tevatron

  • Tevatron status as of September 21

SwitchYard

Continuing with maintenance work

NuMI

  • Complete HV101-1 magnet repair
    • Installed replacement magnet & aligned
    • Vacuum hookup, pumpdown, & certification, hi-pot good
    • Ceramic insulator found broken at water turn on
    • TD repair of ceramic insulator completed in MI tunnel
    • Good Hi-Pot of complete tunnel system & cables to PS
    • Final survey ongoing
  • Target, Horn 1 & Horn 2 maintenance and upgrades
    • Turn on Target/Horn 1 LCW
    • Re-work new Target LCW level transmitter
    • Weld three weldolets and all flanges of the Horn 2 water lines
    • Install three pressure gauges for the Horn 2 supply lines
    • Install close-loop lines for 5 hr test run of Horn 2, check for leaks

MiniBooNE

No work was completed this last week.

  • Reverse Horn Polarity
    • Take wipes. (week of Sep. 10)
    • Reverse polarity of horn power supply. (week of Sep. 17)
      • Test horn
  • Check Corrosion Tabs at 25m Absorber (27/28Aug07)
  • Change LCW DI Bottle (50%)
  • Dry out horn (09Aug07)
  • Check Leakage Current on Ion Pumps
  • Power Supply Maintenance
  • Toroid Work

MuCool Beamline Installation

  • Electrical Work (behind schedule)
    • Cable pull complete
      • Termination of load cables ongoing (50%)
      • Termination of the instrumentation cables ongoing (gallery and tunnel)
    • Interlock box installed and electrically terminated
    • Next week: Finalize cable termination and instrumentation connections
  • Alignment
    • Beam line has been rough aligned
      • Some issues were revealed:
      • Magnets stationing of most components were unsatisfactory
        • Quadrupole magnets were repositioned and realigned
          • Cause: installation error
        • CDA magnets still need to be repositioned and realigned
          • A new nominal values were released for these magnets
      • Beam sleeve thru the shielding wall has been aligned
      • AMG started referencing the MW cans.
    • Next week: - Aid on the adjustment of the CDA magnets station
      • Continue referencing the MW cans
  • LCW work
    • System was hydrostatic pressure tested
    • Water techs are fabricating hoses and fittings
    • Next week: Start making hose-connection between the manifold and magnets
  • Vacuum work
    • Spool piece progress was delayed due to problems with magnets stationing
      • Quadrupoles spool pieces resumed middle of the week
    • MW cans assembly ongoing
    • Next week:  Finalize the spool piece connections
      • Receive and install the MW vacuum cans
      • Start roughing vacuum by the end of the week
  • Mechanical Installation
    • Beam stop mechanical parts have been delivered from the machine shop
      • Currently parts have been cleaned
      • Final assembly started
    • Next week: Install the beam stop (upstream)
  • Shielding installation
    • Shielding added around the beam pipe sleeve
    • Gathering material to complete the shielding requirements in the utility penetration
    • Next week: Finalize the shielding requirements for beam operation in Linac

Collider Detector at Fermilab

  • Completed repair work on ISL/L00 cooling
    • Saturday: last epoxy deposited on west Portcard manifolds
    • Monday: Successful vacuum tests of east and west ISL Portcard circuits
    • Tuesday: Flow re-established to East ISL/L00 Portcards for first time since May 13.   Nominal flow and pressures
    • Flow to west in about 1 week (epoxy cure)
    • Thursday: closed east end-plug.   Started checkout of ISL and SVX
  • Portcard Status
    • East Portcard:
      • Deposited epoxy on all 8 manifold/oval tube junctions
      • Leak rate (measured with air pressure test) reduced by factor of about 1500
      • Definitely good enough to operate the ISL and L00
      • Left open for possible final try when done with west
    • West Portcard:
      • Deposited epoxy on all 8 manifold/oval tube junctions
      • Leak rate reduced by fact of about 400, currently about a factor of 3-4 worse than east
      • Definitely good enough to operate the ISL and L00
      • Will put in 1-2 more days to reduce the leak rate
  • Ladder Lines
    • This morning started inspecting east ladder cooling lines
    • Epoxy looks good
    • No pitting observed in Aluminum
    • Stuck prism has detached from the wire and moved deeper into the line
      • Will try to retrieve the wire this afternoon and look for prism
    • No plan to do any repair on these lines based on the inspection
  • Next week:
    • Close east end-plug
    • Detector elevation survey and low beta quad alignment

D-Zero

Work continues

Alignment

Completed Jobs

  • Tevatron Task Manager: Bocean 
    • 5494 As found low beta quads and separators at B0   Date: 8/06/2007- 8/08/2007 Complete
    • 5495 As found low beta quads and separators at D0. Date: 8/09/2007- 8/14/2007 Complete
    • 5498  Measure rolls on 111 dipoles and quads in Tevatron (100 +) Complete
    • 5501 Un roll D16-1 quad when D-1 is warmed.(As Found)  Date: 8/13/2007 Complete
    • 5591 As Founds of the Roman pots at D0   4 of 6 Date: 8/13/2007 85% Complete
    • 5598 A0 dump as found (no request for this job)  Date: 8/23/2007 Complete
    • 5594 As found Schottkies, Separators and Beam pipe at A17 s Date: 8/24/2007   Complete
    • 5572 As found and re survey kicker replacement at F-17.  Date: 8/28/2007   Complete
    • 5506  As found crystal collimator at E0.  Date: 8/28/2007    Complete
    • 5506  level run through Tevatron tie rods.  Date: 8/29/2007  Complete
    • As found TEL2 in transfer hall.              Date: 9/18/2007   Complete
    • 5599  As found E 12-2 12-3 and 12-4 in preparation to pull E 12-3
    • 5499 Measure and re align 29 dipoles and quads that have new stands Date: 9/04/2007- 9/07/2007
    •  As found TEL1 in transfer hall.   Date: 9/17/2007 Completed
    • 5603  As found and set new spool at E 12-1.  Date: 9/17/2007    Completed
  • Main Injector Task Manager: Oshinowo
    • 5563  AS found survey of Quadrupoles at 400, 402, 404 and the three Lambertson. Date: 8/13/2007- 8/15/2007 Complete
    •  5554   307, 308 Collimators: Base plate survey  Date: 8/16/2007 Complete
    • 5554   307, 308 Collimators: Set travel and limit switches Date: 8/22/2007 Complete
    • 5596 Magnet 614 had to be moved please realign. Date: 8/23/2007 Complete
    • 5562 The beam pipe on the downstream side of the MI40 Lambertson.    Date 8/29/2007 Complete
    • 5553  Align new kicker magnet in Main Injector at location Q103.  Date 9/17/2007 Complete
  • Nova Task Manager: Oshinowo
    • 5475- Stake out new Recycler injection line (MI10) Date: 8/27/2007 Complete
  • Booster  Task Manager: Oshinowo
    • 5534 Align 24 Booster Corrector Magnets.( Long 19   Long 17 )   Date: 8/13/2007    Complete
    • 5534 Align 24 Booster Corrector Magnets  ( Long 21   Long 23 ) Date: 8/14/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets  ( Long13     Long15 ) Date: 8/20/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets  ( Long 11   Long 9 ) Date: 8/21/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets  ( Long 7   Long 5 ) Date: 8/23/2007 Complete
    • 5534 Dog Leg BDLA005           Date: 9/18/2007 Complete
    • 5535 As-Set Q800 SQA510.     Date: 9/20/2007 Complete
  • 400 MeV Task Manager: Oshinowo 
    • 5538 Possibility to survey few 400MeV component(s) (2 quads?). Date: 9/19/2007
  • P-Bar Task Manager: Oshinowo
    • 5488 As Found A4Q10 (LQA in Accumulator) has a water leak. Date: 8/09/2007 Complete
    •  5595 As Found An Accumulator multipole magnet has developed a leak. Date: 8/30/2007 Complete
    • 5488 Alignment A4Q10 (LQA in Accumulator) has a water leak.   Date: 8/30/2007 Complete
    • 5595 Alignment An Accumulator multipole magnet has developed a leak. Date: 9/04/2007 Complete
    • 5540 The Debuncher DCCT (located between D1Q2 and D1Q3). Date: 9/12/2007 Complete
    • 5541  A new three-piece ceramic tube for the Debuncher Extraction Kicker (located between D6Q10 and D6Q9) will be installed. Date: 9/18/2007 Complete
  • Recycler Task Manager: Oshinowo
    • 5527  Optically align Recycler DCCT after installation. Date: 8/13/2007 Complete
    • 5600  AS found all beam line components in the R32 transfer line between S803A and S804B. Complete
    • 5601 Multiwire MW803 in the Recycler 800 line had a leaky seal. Date: 9/11/2007 Complete
  • Linac Task Manager: Oshinowo
    • 5468  Stake MuCool beamline component locations in the Linac and (MTA). Date: 8/08/2007 Complete
  • CMS Task Manager: Oshinowo
    • 5516 Survey the CMS F-pixel Half cylinder +Z2 in the Lab C Date: 8/09/2007 Complete
    • 5518 Survey the CMS F-pixel Half cylinder +Z2 in the Lab C    Date: 9/06/2007 Complete
  • D0 Task Manager: Oshinowo
    • Stickmike survey of Muon system prior to opening of detector. This will include both CFs and EFs. Date: 8/06/2007  Complete
    • 5524  VStar or Laser Tracker survey of North/South EF pedestals with respect to the sidewalk monuments. Date: 8/07/2007    Complete
    • 5592 Elevation run in D0 before machine is opened. Date: 8/06/2007 Complete
  • P150 Task Manager: Oshinowo  Late Request
    • 5593  We had to move several magnets and multiwires to make up the ends on the cans for vacuum connections.  They are as follows:  The multiwire cans at 708 and 710.  B2 magnets at the interface of 710 - 711 and the interfaces of 709 - 710, the serial #'s are BDM 883-1  BDM 835-1 BDM 879-1 and BDM 397-1. Date: 9/07/2007 Complete
  • NuMI Task Manager: Bocean
    • 5602  As found the currently existing HV101-1 (spare EDBB017). Date: 9/11/2007  Complete
    • 5602  As Set the currently existing HV101-1 (spare EDBB017). Date: 9/19/2007  Complete
  • MTA Task Manager: Oshinowo
    • 5469 Vertical check of the MuCool beam line components. Date: 8/29/2007 Complete
    • Shoot MTA shielding blocks after re-stacking. Date: 9/04/2007 Complete
    • 5469  Rough alignment of the MuCool beam line components. Date: 9/20/2007 ?

Next week’s schedule of work

  • Booster   Task Manager: Oshinowo
    • Align new kickers (the new corrector will be aligned at the same time) at long 2 after installation
  • Main Injector Task Manager: Oshinowo
    • 5554  4 collimators in the Main Injector at the 301, 303, 307, and 308 locations. Date:9/20/2007- 9/24/2007
  • Elevation Run Floor points Manager: Kyle
    • 5528  Ring Wide elevation survey            Date: 9/17/2007
  • Tevatron Task Manager: Bocean
    • 5572  As found and re survey kicker replacement at F-17. Scott McCormick will notify when ready. Date: ?
    • 5496  Re align low beta quads and separators at B0    Date:9/27/2007
    • 5497  Re align low beta quads and separators at D0    Date:9/27/2007
  • Nova Task Manager: Oshinowo 
    • 5477-Stake out new Recycler RR30 quad positions for the Nova Project  Date: 9/20/2007
    • 5476-Stake out new Recycler extraction line (MI30) for the Nova Project Accelerator upgrade     Date: 9/20/2007
  • CDF Task Manager: Greenwood
    • 5450  Measure elevation of CDF central detector Elevation Run. Date: 9/24/2007
  • NuMI Task Manager: Bocean
    • 5602 Align the HV101-1 magnet (the initial dipole EDBB020) after re-installation at MI-60. Date: 9/15/2007
  • ILC Task Manager: Bocean Not Scheduled

Alignment and Metrology Home page

Mechanical

Work in the tunnels continues.

Cryogenic System

Tevatron at LN2 temperature except A1 and E1 at room temperature

  • Pbar Debuncher Stochastic tanks and S/C Notch Filter Dewar are at LHe temperature
  • Tunnel magnet cryostat vacuums at A25 and D35 regions are of concern.
  • 3 FRIG Mycom compressors (plus CDF & DZero) running with helium precooled to LN2 temperature flowing through magnet strings
  • CHL helium inventory management system operating on clean Helium Tank Farm,  though A0 GHe Recovery being vented due Recovery compressor being replaced.  CHL Main Helium Purifier was derimed, cooled down, and put back online to Screw Compressor.  CHL Mobile Helium Purifier was taken offline and will be derimed at room temperature
  • CHL Nitrogen Reliquefier (NRL) plant off and LN2 supply dependent on vendor deliveries to CHL LN2 Dewar tanks; deliveries are fine.  NRL IR Booster compressor startup was attempted but took down north-end bldg power
  • CHL 24/7 operations shiftwork resumes Oct. 1
  • Repair options for feeder 38 cable between CHL and FRIG B0/BA Compressor that failed in August need discussion between FESS, CDF, and Cryo

FRIG Maintenance

  • 13 of 13 wet expander flywheels have been replaced
  • 25 of 26 wet/dry expander overhauls of cold ends are completed (only E1 left to do after tunnel vacuum work is completed)
  • Various vacuum, instrumentation, and controls maintenance continuing.
  • Mycom compressor startup checkout in progress

CHL Maintenance

  • Coldbox-1:
    • Charcoal Adsorber west unit regeneration was completed and both east and west units will be purified next week
  • Nitrogen Reliquefier plant maintenance in progress:
    • Marley Cooling tower was test run fine after overhaul
    • IR Booster compressor power issues being investigated
  • Various compressors and facility maintenance in progress

Controls

This Week

  • TLG Reboot
  • Database Disk Array Firmware Patch Installations – Complete
  • Unix Webserver Maintenance – Complete
  • Save/Restore and SDA Database Updates – Complete
  • Alarms Server port to Linux - Complete
  • NuMI AAT PLC Upgrade – Complete
  • Linac Tank 1 Collimator Controls Upgrade
  • MI-30 Collimator Controls Installation 
  • Monitored Sinkhole Repair at E3 (19 Conductor Proximity)
  • Routine maintenance

Next Week

  • 9/22 & 23/07
    • No Control System, No Accelerator Clock Systems
      • From ~7:00 AM Saturday through Sunday Evening
  • 9/27/07
    • Application and Device Database Updates

FESS

CUB Shutdown 9/22 – 9/23

  • FDR 42 – Servicing 13.8KV switches from Giese Rd to CUB
  • Servicing 13.8kv/4160v – 480v unit sub-stations/transformers for CUB
  • Servicing electrical 480/120v panels/switches inside CUB
  • DDC work – control wiring/upgrades/system checks
  • Cooling tower sump/tank check
  • Sunday – complete water system checks & start-up all equipment/cooling systems by 5:00pm

Casey’s Pond

  • OPERATING on the 400HP PUMPS [flowing thru 14 inch strainer]
  • C-4 ICW Pumps not in operation unless required
  • ICW FLOW METER WIP.  Expected completion/transfer back to the 16 inch strainer by 9/28/07
  • Two intake pipes cleaned out – 6,000lbs of ZM removed
  • New ICW liquid Bromine injection treatment to start next week

Poles

  • ALL 0f the19 poles are installed with conductors attached.
  •  Testing New Breaker/Air Switches next week
  •  Power Up lines by 9/29/07 into MSS yard prior to site outage/MSS energizing on 10/3/07

TeV pond silt and Extension Collars

  • Silt transfer completed for all MR sectors
  • Suction screen extensions WIP – 10 extensions installed & 11 more scheduled.  Expect to complete during SD

Other Jobs

  • CUB – 55LCW PUMP REPLACEMENTS WIP.  LP 4 is completed & scheduled to run today.  Should be able to keep flow while LP5 is replaced.  May require 55 LCW short shutdowns to finalize project
  • TENTATIVE PROCESS CHILLED WATER SHUTDOWN TO Pbar, MI & NUMI ON WEDS 9/26/07 for removal of flow control device at CUB for 2 – 4 hours that is impeding flow
  • TENTATIVE SITE ICW FLUSH SCHEDULED TO START SAT 9/29/07 THRU 10/06/07
  • ZEBRA MUSSEL TREATMENT SCHEDULED FOR MID-OCTOBER

Safety

  • Safety System Status
    • Upgrade M02 - M05 Enclosure Interlocks                                            80%
    • Upgrade M01 Enclosure Interlocks                                                      55%
    • Upgrade Main Injector PLC Code and replace PLC Batteries              50%
    • Upgrade Meson Area Oxygen Monitors                                                10%
    • Modify Linac Enclosure Interlocks                                                     100%
    • Replace Safety System DEC talk with ATT Natural Voices                30%
    • Replace Pbar Vault Key Tree Door Latches & add status LEDs             0%
    • Linux kernal 2.4 to 2.6 upgrade to Safety Data Acquisition System   100%
    • Route Key Tree release signals from MCR to Meson                          100%
    • Install Radiation Monitor Cables for the Muon Test Area                   100%
    • Replace oxygen monitor cells in the Tevatron                                     100%
  • Startup
    • Linac Safety System requires testing before startup
    • Machine Startup Approval Documents available from Mike Gerardi
  • A successful shutdown is getting our work done safely

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/21/2007    email Fermilab